# Tomur: Traffic-Aware Performance Prediction of On-NIC Network Functions with Multi-Resource Contention

### v3

@@ 2.3 Challenge 2: Traffic Attributes
Figure 3: (a) FlowStats’s throughput when the competitor’s cache access rate (CAR) changes in three distinct traffic profiles. CAR is the sum of the cache read and write rates obtained from the hardware performance counters on BlueField-2. (b) Distribution of prediction errors after adapting the model to different traffic profiles. We show the median error on top of each error box.
3 Tomur Overview
The challenges in 2 pose two fundamental questions on accurate performance prediction for on-NIC NFs, which drive Tomur’s design:
  1. 1.
@@ 3 Tomur Overview
We develop Tomur, a framework for accurately predicting on-NIC NF performance with multi-resource contention and varying traffic profiles. To address the first design question, Tomur adopts a “divide-and-compose” approach: it builds up individual per-resource contention models (4.1) for both hardware accelerators and memory subsystem to separately model their impact on throughput, and applies execution-pattern-based composition (4.2) to faithfully capture the end-to-end effect of contention. Then in response to the second design question, Tomur introduces traffic-aware augmentation (5.1) techniques to integrate various traffic attributes into per-resource models, and develops an adaptive profiling (5.2) method to balance the soaring profiling costs (due to the extra dimensions of traffic attributes) with model quality. Taken together, during online prediction, Tomur takes the contention level of competing NFs and traffic attributes of the target NF as input to the per-resource models and compose the results based on NF’s execution pattern to obtain the final prediction. Detailed workflow is in Appendix F. Consistent with prior work [42, 44, 32], Tomur does not require knowledge of or access to NF source code.
4 Multi-Resource Contention Modeling
@@
We now present the design insights and details of Tomur’s multi-resource contention modeling. Note we are interested in the NF’s maximum throughput assuming the arrival rate is high enough, which represents the NF’s capability and is consistent with prior work [42, 44].
4.1 Per-Resource Models
@@
An on-NIC NF consumes onboard CPU, memory subsystem (cache and main memory), hardware accelerators, and NIC [49, 31, 35, 32]. For CPU, given common deployment practice [56, 54, 60], we perform core-level isolation for co-located NFs so CPU contention does not happen. Although some prior work has discussed potential isolation issues for NICs on a server [35], we do not encounter this problem as on-NIC NFs leverage powerful hardware-based flow table on SmartNICs [6]. Thus, we focus on contention on hardware accelerators and memory subsystem here, assuming fixed traffic attributes. Notice the per-resource modeling effectively derives the NF’s throughput on one given resource only without accounting for other resources, and may not equal to the overall throughput.
4.1.1 Hardware Accelerators
@@
At first glance, modeling hardware accelerator contention seems not much different from existing design for memory contention [42, 44]. That is, one can use an accelerator’s performance counters to quantify NF’s contention level as the input, and employ an ML model to predict throughput. This is infeasible, however, because current SmartNIC accelerators do not expose fine-grained performance counters [10, 3, 39, 19, 12]. We thus propose a general queue-based white-box approach for hardware accelerators.
@@
Contention behavior in hardware accelerators. We start by analyzing the accelerator’s contention behavior which our modeling is based upon. Without loss of generality, we use the widely-used regex accelerator [20, 6, 12] as the target of discussion hereafter.
@@
In practice, NFs utilize onboard accelerators via the corresponding queue systems [7, 6]. For example, an NF establishes request queues and enqueues/dequeues operations to/from a regex accelerator [7, 20, 6]. This queue-based interface unifies the interaction with specialized accelerators and applies to many SmartNICs [13, 10, 11] and beyond [33, 1]. Understanding the queue system behavior is then crucial for modeling accelerator contention.
@@
These two observations are very familiar to us as they point to the canonical round-robin (RR) queuing discipline widely used in practice. Indeed, we confirm from [9] that our regex accelerator driver’s implementation adopts RR for queue-level fairness. With one queue per each NF which is the setup in our experiment, as regex-bench’s request arrival rate increases, regex-NF’s requests have proportionally less access to the accelerator, resulting in the linear throughput decline. When contention is high enough that regex-bench’s queue is always non-empty when the RR scheduler turns to it, throughput of regex-NF stops dropping since its average sojourn time (sum of queueing and processing times) stops increasing [46]. As they have the same numbers of queues, their equilibrium throughput is the same as seen in Figure 4.
Figure 4: Throughput of co-running synthetic pattern-matching “regex-NF” and regex-bench as a function of arrival rate of regex-bench. The packet size is 1500B. In each setting, regex-NF and regex-bench reach an equilibrium throughput, e.g., , with MTBR of 194 maches/MB for regex-NF, they both obtain 48.9 Mpps at equilibrium.
Our approach. Motivated by the above analysis, the regex accelerator contention can be modeled by RR over multiple request queues with one service node (i.e., the accelerator). Suppose we have $N$ NFs sharing an accelerator, and each $NF_{j}$ has $n_{j}$ request queues. At equilibrium, the average sojourn time $t$ of requests from each queue is [46]:
$t$ $=\sum_{j=1}^{N}n_{j}t_{j},$ (1)
where $t_{j}$ represents $NF_{j}$’s average request processing time. For a target $NF_{i}$, its throughput $T_{i}$ can be represented as the sum of throughput of all its queues, i.e.,
$T_{i}$ $=\frac{n_{i}}{t}=\frac{n_{i}}{\displaystyle\sum_{j=1}^{N}n_{j}t_{j}}=\frac{n_{i}}{\displaystyle\sum_{j=1}^{N}\frac{n_{j}^{2}}{T_{j,solo}}},$ (2)
@@
We leverage Equation (2) to derive the equilibrium throughput under two one-queue NFs (Appendix E) just as in Figure 4, and verify with empirical results of various regex-based NFs (Appendix A) that our approach is accurate with 1.3% error on average.
Other accelerators. Our approach here directly applies to other hardware accelerators, e.g., compression and crypto accelerator, which also uses round-robin based queues [9, 17].
4.1.2 Memory Subsystem
Memory subsystem contention has been studied in existing work [42, 44, 28] which finds that the contention-induced throughput drop can be modeled as a piece-wise linear function of performance counters (Figure 9 in Appendix B). Thus we follow SLOMO’s gradient boosting regression (GBR) method which is state-of-the-art, using 7 performance counters (Table 13 in Appendix D) as input features. Note that we overcome the fixed-traffic limitation of GBR by integrating traffic attributes to it in 5.
4.2 Execution-Pattern-Based Composition
@@
Observations. We analyze two typical execution patterns of NFs: pipeline and run-to-completion [32, 20]. Considering a packet received by a pipeline NF, or p-NF, and a run-to-completion NF, or r-NF: for p-NF, the packet waits at the first stage until its predecessor enters the second stage; for r-NF, the packet waits until the predecessor leaves the last stage.
@@
$T$ $=T_{solo}-\max(\Delta T_{1},...,\Delta T_{r}),$ (3)
@@
$T$ $=\frac{1}{\displaystyle\sum_{k=1}^{r}(t_{k}+\Delta t_{k})}$ (4)
$=\frac{1}{\displaystyle\sum_{j=1}^{r}\left(t_{j}+\Delta t_{j}+\sum_{k=1,k\neq j}^{r}t_{k}\right)-\sum_{j=1}^{r}\sum_{k=1,k\neq j}^{r}t_{k}}$
@@
Detecting execution pattern. Without source code access, we resort to a simple testing procedure to detect an NF’s execution pattern. We co-run the NF with our benchmark NFs, and see if Equation 3 or 4 fits its throughput drop better. One may also observe the NF’s throughput curve similar to Figure 5 to empirically determine if it is a p- or r-NF.
Figure 5: Throughput of two synthetic Click NFs that use pipeline (top) and run-to-completion (bottom) as a function of competing CAR in memory subsystem and match rate in regex accelerator.
5 Traffic-Aware Prediction
@@
5.1 Traffic-Aware Augmentation
It is obvious that we need to augment the per-resource model with knowledge of traffic attributes, while execution-pattern-based composition is not affected. To do this, we select three common traffic attributes that impact NF performance based on our experiment results and previous studies [42, 32]: number of flows or flow count, packet size, and match-to-byte-ratio (MTBR) of a packet. We denote a traffic profile of 16K flows, 1500B packets and 600 matches/MB payload using a vector $(16000,1500,600)$.
5.1.1 Hardware Accelerators
@@
Our approach. A regex-utilizing NF is naturally sensitive to the MTBR of packet payload [20]. [^4] Number of queues is fixed during NF’s life cycle per its configuration. It directly impacts the average processing time of the regex requests.
Following notations from Equation 2, the average request processing time of $NF_{j}$ for a given ruleset can be expressed as:
$t_{j}=\frac{1}{T_{j,solo}}=t_{j,0}+a_{j}m_{j},$ (5)
where $t_{j,0}$ and $a_{j}$ are constants, and $m_{j}$ is the MTBR. Equation 5 builds on the intuition that that request processing time grows linearly with number of matches in a packet, as each match induces a constant amount of extra processing time on average on top of the base processing time. Plugging into Equation 2, the traffic-aware throughput model of regex accelerator is:
$T_{i}$ $=\frac{n_{i}}{\sum_{j=1}^{N}n_{j}^{2}(t_{j,0}+a_{j}m_{j})}.$ (6)
@@ 5.1.2 Memory Subsystem
Figure 6: Throughput of FlowStats as a function of traffic attributes. (a) The packet size is 1500 B. (b) The number of flows is 16 K.
@@
Observation. We co-run FlowStats with mem-bench with different numbers of flows (other traffic attributes stay the same). We adjust the working set size of mem-bench and keep other metrics (i.e., CAR and memory access rate) fixed. Figure 6(a) reveals that FlowStats experiences significant throughput drop with increasing flow count. More interestingly, the throughput curve is also a piece-wise function of the number of flows, which is similar to memory contention modeling as mentioned in 4.1.2.
Our approach. Thus, we take a straight-forward approach to add traffic attribute knowledge as extra features to the blackbox GBR model of memory subsystem, by appending the traffic attributes vector to the original input vector of performance counters.
5.2 Adaptive Profiling
@@
Being traffic-aware brings a new and critical issue: training a traffic-aware model now needs more data due to the higher dimensionality of the input vector compared to a fixed-traffic model. This is especially true for the black-box model for memory subsystem, which already has a high-dimensional input vector. A naive solution is to repeat the collection process for each unique traffic profile (a.k.a. full profiling), which obviously leads to massive exponentially-growing profiling overheads. Although random sampling based profiling [42, 44] can effectively reduce overhead, it may cause high prediction error due to inadequate coverage of the data space. The trade-off between profiling cost and model accuracy motivates us to design an optimized profiling method for Tomur.
@@
Observations. Our design has roots on two key observations. First, an NF’s performance is only dependent on a few traffic attributes. For example, FlowStats is only sensitive to number of flows as in Figure 6(a), but not packet size as in Figure 6(b), since it only processes packet header information.
@@
Second, for a given traffic attribute, it causes significant performance drops only in a limited range of values; performance shows little changes in other ranges. Still using FlowStats, when the competing WSS is 10MB, its throughput loses $\sim$30for $[1,20K]$ flows, but remains almost unchanged for $[40,60K]$ flows as in Figure 6(a). This is because FlowStats’s hash table grows as a result of growing flow count. Before it fully occupies the last level cache (LLC), increasing the hash table size leads to higher cache miss ratio, which slows down read/write operations and thus throughput. After the LLC is saturated, cache miss ratio stays at a fixed level, causing NFs to exhibit constant throughput. These characteristics also hold in general across NFs as we empirically observe for other NFs in Table 1. This is because generally, (1) NFs typically process either packet headers or payloads, which only depends on several traffic attributes; and (2) traffic attributes usually affect performance by changing the size of key data structures in the NF processing logic, e.g., the mapping table in NAT [42, 29, 44], which exhibits the same LLC effect as explained just now.
@@
As shown in Algorithm 1, we first test whether NF performance is sensitive to a traffic attribute or not. Suppose the possible range of an attribute $f$ is $[f_{min},f_{max}]$. We profile the NF’s solo throughput using with $f$ set to $f_{min}$ and $f_{max}$ respectively while other attributes remain default (lines 8-9). Then we compare the throughput difference to determine if $f$ should be added in our model. Note that the function profile_one() collects throughput data under a specified configuration (contention level and traffic attribute), and increments the total number of collected samples by one if the configuration has not been profiled. After pruning the attribute list, we carry out a binary search to adaptively collect profiling data. For each call to RANGE_PROFILE(), we consider the difference between solo throughputs on traffic attribute boundaries. If their difference exceeds a certain threshold, we collect $m$ data points at the center of current traffic attribute region (lines 18-22), Then the traffic attribute region is split in half, which will serve as the new region for the recursive call of RANGE_PROFILE().
@@
2:$C$: The list of performance counters
3:$F$: The list of traffic attributes
@@
5:function adaptive_profile($NF$, $C$, $F$)
@@
7:    for $f$ in $F$ do
8:         $T_{min}$leftarrow profile_one($NF$, 0, $f_{min}$, $n$) $\triangleright$ "0" represents no contention.
9:         $T_{max}$leftarrow profile_one($NF$, 0, $f_{max}$, $n$)
@@
11:             $F$leftarrow $F$$-f$triangleright Prune the traffic attribute dimension.              RANGE_PROFILE($NF$, $F_{min}$, $F_{max}$, $n$)
12:    return
13:function range_profile($NF$, $F_{min}$, $F_{max}$, $n$)
14:    $T_{min}$leftarrow profile_one($NF$, 0, $F_{min}$, $n$)
15:    $T_{max}$leftarrow profile_one($NF$, 0, $F_{max}$, $n$)
16:    if $n$geq $Q$ then $\triangleright$ We reach profiling quota
17:         return
18:    if $|T_{max}-T_{min}|$geq $\epsilon_{1}$ then $\triangleright$ Only enforce collection when throughput changes drastically.
19:         $F_{mid}$leftarrow $\frac{F_{max}+F_{min}}{2}$
20:         for _ in $m$ do
21:             $C_{r}$leftarrow random() $\triangleright$ Contention level is random.
22:             profile_one($NF$, $C_{r}$, $F_{mid}$, $n$)
23:         RANGE_PROFILE($NF$, $F_{mid}$, $F_{max}$, $n$)
24:         RANGE_PROFILE($NF$, $F_{min}$, $F_{mid}$, $n$)
25:    return
6 Implementation
We implement Tomur with $\sim$1600 LoCs in C and Python. Tomur employs perf-tools [19] to collect performance counters (appendix D). Tomur uses sklearn [21] to construct machine learning models used in per-resource models. Our code is open source anonymously at [14].
Synthetic benchmarking NFs. We implement three synthetic NFs called mem-bench, regex-bench and compression-bench ($\sim$8300 LoCs in C with DPDK support) based on open-source benchmark tools [22, 18, 20, 7] to apply configurable levels of contention on memory subsystem, regex, and compression accelerators, respectively. The mem-bench accesses an array the system memory by varying: (1) access pattern, i.e., stream access or step access, (2) access speed, and (3) the size of the array. regex-bench can adjust its processing rate and input profiles to induce different request rates and per-request Match-to-Byte Ratios (MTBR). Similarly compression-bench also varies its processing rate to induce different request rates and contention on the compression accelerator.
Network functions. We implement common on-NIC NFs with $\sim$3600 LoCs in C and Click using frameworks including Click 2.1 [29], DPDK 20.11.6 [17], and DOCA 1.5-LTS [6]. Our BF-2 enables hardware flow table offloading for NFs. We present how we use these NFs to evaluate Tomur in 7.1.
7 Evaluation
We present our evaluation of Tomur now. The highlights are:

### v5

@@ 2.2.2. Traffic Attributes
Figure 3. (a) FlowStats’s throughput when the competitor’s cache access rate (CAR) changes in three distinct traffic profiles. CAR is the sum of the cache read and write rates obtained from the hardware performance counters on BlueField-2. (b) Distribution of prediction errors after adapting the model to different traffic profiles. We show the median error on top of each error box.
[…]
3. Yala Overview
The challenges in 2 pose two fundamental questions on accurate performance prediction for on-NIC NFs, which drive Yala’s design:
  1. (1)
@@ 3. Yala Overview
We develop Yala, a framework for accurately predicting on-NIC NF performance with multi-resource contention and varying traffic profiles. To address the first design question, Yala adopts a “divide-and-compose” approach: it builds up individual per-resource contention models (4.1) for both hardware accelerators and memory subsystem to separately model their impact on throughput, and applies execution-pattern-based composition (4.2) to faithfully capture the end-to-end effect of contention. Then in response to the second design question, Yala introduces traffic-aware augmentation (5.1) techniques to integrate various traffic attributes into per-resource models, and develops an adaptive profiling (5.2) method to balance the soaring profiling costs (due to the extra dimensions of traffic attributes) with model quality. Taken together, during online prediction, Yala takes the contention level of competing NFs and traffic attributes of the target NF as input to the per-resource models and compose the results based on NF’s execution pattern to obtain the final prediction. Consistent with prior work (slomo20, ; bubbleup11, ; lognic23, ), Yala does not require knowledge of or access to NF source code.
4. Multi-Resource Contention Modeling
@@
We now present the design insights and details of Yala’s multi-resource contention modeling. Note we are interested in the NF’s maximum throughput assuming the arrival rate is high enough, which represents the NF’s capability and is consistent with prior work (slomo20, ; bubbleup11, ; clara21, ).
4.1. Per-Resource Models
@@
An on-NIC NF consumes onboard CPU, memory subsystem (cache and main memory), hardware accelerators, and NIC (clara21, ; fairnic20, ; picnic20, ; lognic23, ). For CPU, given common deployment practice (resq18, ; nfp17, ; opennetvm16, ), we perform core-level isolation for co-located NFs so CPU contention does not happen. Although some prior work has discussed potential isolation issues for NICs on a server (picnic20, ), we do not encounter this problem as on-NIC NFs leverage powerful hardware-based flow table on SmartNICs (ovsdpdk, ). Thus, we focus on contention on hardware accelerators and memory subsystem here, assuming fixed traffic attributes. Notice the per-resource modeling effectively derives the NF’s throughput on one given resource only without accounting for other resources, and may not equal to the overall throughput.
4.1.1. Hardware Accelerators
@@
At first glance, modeling hardware accelerator contention seems not much different from existing design for memory contention (slomo20, ; bubbleup11, ). That is, one can use an accelerator’s performance counters to quantify NF’s contention level as the input, and employ an ML model to predict throughput. This is infeasible, however, because current SmartNIC accelerators do not expose fine-grained performance counters (bf21, ; regexbf, ; bf2perf21, ; perf, ; bf2forum, ). We thus propose a general queue-based white-box approach for hardware accelerators.
@@
Contention behavior in hardware accelerators. We start by analyzing the accelerator’s contention behavior which our modeling is based upon. Without loss of generality, we use the widely-used regex accelerator (RXPbench, ; doca1, ; bf2forum, ) as the target of discussion hereafter.
@@
In practice, NFs utilize onboard accelerators via the corresponding queue systems (DPDK, ; doca1, ). For example, an NF establishes request queues and enqueues/dequeues operations to/from a regex accelerator (DPDK, ; RXPbench, ; doca1, ). This queue-based interface unifies the interaction with specialized accelerators and applies to many SmartNICs (pensando, ; bf21, ; bf31, ) and beyond (KKA16, ; rdma-qp, ). Understanding the queue system behavior is then crucial for modeling accelerator contention.
@@
These two observations are very familiar to us as they point to the canonical round-robin (RR) queuing discipline widely used in practice. Indeed, we confirm from (mlxregex, ) that our regex accelerator driver’s implementation adopts RR for queue-level fairness. With one queue per each NF which is the setup in our experiment, as regex-bench’s request arrival rate increases, regex-NF’s requests have proportionally less access to the accelerator, resulting in the linear throughput decline. When contention is high enough that regex-bench’s queue is always non-empty when the RR scheduler turns to it, throughput of regex-NF stops dropping since its average sojourn time (sum of queuing and processing times) stops increasing (queue1, ). As they have the same numbers of queues, their equilibrium throughput is the same as seen in Figure 4.
Figure 4. Throughput of co-running synthetic pattern-matching “regex-NF” and regex-bench as a function of arrival rate of regex-bench. In each setting, regex-NF and regex-bench reach an equilibrium throughput, e.g., , with MTBR of 194 maches/MB for regex-NF, they both obtain 48.9 Mpps at equilibrium.
[…]
Our approach. Motivated by the above analysis, the regex accelerator contention can be modeled by RR over multiple request queues with one service node (i.e., the accelerator). Suppose we have $N$ NFs sharing an accelerator, and each $NF_{j}$ has $n_{j}$ request queues. At equilibrium, the average sojourn time $t$ of requests from each queue is (queue1, ): $t=\sum_{j=1}^{N}n_{j}t_{j}$, where $t_{j}$ represents $NF_{j}$’s average request processing time. For a target $NF_{i}$, its throughput $T_{i}$ can be represented as the sum of throughput of all its queues, i.e.,
(1) $T_{i}$ $=\frac{n_{i}}{t}=\frac{n_{i}}{\displaystyle\sum_{j=1}^{N}n_{j}t_{j}}=\frac{n_{i}}{\displaystyle\sum_{j=1}^{N}\frac{n_{j}^{2}}{T_{j,solo}}},$
@@
We verify Equation (1) with empirical results of various regex-based NFs, which show that our approach is accurate with 1.3% error on average.
Other accelerators. Our approach here directly applies to other hardware accelerators, e.g., compression and crypto accelerator, which also uses round-robin based queues(comprbf, ; pkabf, ).
4.1.2. Memory Subsystem
Memory subsystem contention has been studied in existing work (slomo20, ; bubbleup11, ; dobre12, ) which finds that the contention-induced throughput drop can be modeled as a piece-wise linear function of performance counters. Thus we follow SLOMO’s gradient boosting regression (GBR) method which is state-of-the-art, using 7 performance counters as input features. Note that we overcome the fixed-traffic limitation of GBR by integrating traffic attributes to it in 5.
4.2. Execution-Pattern-Based Composition
@@
Observations. We analyze two typical execution patterns of NFs: pipeline and run-to-completion (lognic23, ; RXPbench, ). In the following discussion, we define a stage as a processing block that will only utilize one resource type. Considering a packet received by a pipeline NF, or p-NF, and a run-to-completion NF, or r-NF: for p-NF, the packet waits at the first stage until its predecessor enters the second stage; for r-NF, the packet waits until the predecessor leaves the last stage.
@@
(2) $T$ $=T_{solo}-\max(\Delta T_{1},...,\Delta T_{r}),$
@@
(3) $T$ $=\frac{1}{\displaystyle\sum_{k=1}^{r}(t_{k}+\Delta t_{k})}=\frac{1}{\displaystyle\sum_{j=1}^{r}\left(t_{j}+\Delta t_{j}+\sum_{k=1,k\neq j}^{r}t_{k}\right)-\sum_{j=1}^{r}\sum_{k=1,k\neq j}^{r}t_{k}}$
@@
Detecting execution pattern. Without source code access, we resort to a simple testing procedure to detect an NF’s execution pattern. We co-run the NF with our benchmark NFs, and see if Equation 2 or 3 fits its throughput drop better. One may also observe the NF’s throughput curve similar to Figure 5 to empirically determine if it is a p- or r-NF.
Figure 5. Throughput of two synthetic Click NFs that use pipeline (top) and run-to-completion (bottom) as a function of competing CAR in memory subsystem and match rate in regex accelerator.
[…]
5. Traffic-Aware Prediction
@@
5.1. Traffic-Aware Augmentation
It is obvious that we need to augment the per-resource model with knowledge of traffic attributes, while execution-pattern-based composition is not affected. To do this, we select three common traffic attributes that impact NF performance based on our experiment results and previous studies (slomo20, ; lognic23, ): number of flows or flow count, packet size, and match-to-byte-ratio (MTBR) of a packet. We denote a traffic profile of 16K flows, 1500B packets and 600 matches/MB payload using a vector $(16000,1500,600)$.
5.1.1. Hardware Accelerators
@@
Our approach. A regex-utilizing NF is naturally sensitive to the MTBR of packet payload (RXPbench, ). [^4] Number of queues is fixed during NF’s life cycle per its configuration. It directly impacts the average processing time of the regex requests.
Following notations from Equation 1, the average request processing time of $NF_{j}$ for a given ruleset can be expressed as: $t_{j}=\frac{1}{T_{j,solo}}=t_{j,0}+a_{j}m_{j}$, where $t_{j,0}$ and $a_{j}$ are constants, and $m_{j}$ is the MTBR. This equation builds on the intuition that that request processing time grows linearly with number of matches in a packet, as each match induces a constant amount of extra processing time on average on top of the base processing time. Plugging into Equation 1, the traffic-aware throughput model of regex accelerator is:
(4) $T_{i}$ $=\frac{n_{i}}{\sum_{j=1}^{N}n_{j}^{2}(t_{j,0}+a_{j}m_{j})}.$
@@ 5.1.2. Memory Subsystem
Figure 6. Throughput of FlowStats as a function of traffic attributes. (a) The packet size is 1500 B. (b) The number of flows is 16 K.
[…]
@@
We co-run FlowStats with mem-bench with different numbers of flows (other traffic attributes stay the same). We adjust the working set size of mem-bench and keep other metrics (i.e., CAR and memory access rate) fixed. Figure 6(a) reveals that FlowStats experiences significant throughput drop with increasing flow count. More interestingly, the throughput curve is also a piece-wise function of the number of flows, which is similar to memory contention modeling as mentioned in 4.1.2 (we shall explain this in 5.2 later). Thus, we take a straight-forward approach to add traffic attribute knowledge as extra features to the blackbox GBR model of memory subsystem, by appending the traffic attributes vector to the original input vector of performance counters.
5.2. Adaptive Profiling
@@
Being traffic-aware brings a new and critical issue: training a traffic-aware model now needs more data due to the higher dimensionality of the input vector compared to a fixed-traffic model. This is especially true for the black-box model for memory subsystem, which already has a high-dimensional input vector. A naive solution is to repeat the collection process for each unique traffic profile (a.k.a. full profiling), which obviously leads to massive exponentially-growing profiling overheads. Although random sampling based profiling (slomo20, ; bubbleup11, ) can effectively reduce overhead, it may cause high prediction error due to inadequate coverage of the data space. The trade-off between profiling cost and model accuracy motivates us to design an optimized profiling method for Yala.
@@
Observations. Our design has roots on two key observations. First, an NF’s performance is only dependent on a few traffic attributes. For example, FlowStats is only sensitive to number of flows as in Figure 6(a), but not packet size as in Figure 6(b), since it only processes packet header.
@@
Second, for a given traffic attribute, it causes significant performance drops only in a limited range of values; performance shows little changes in other ranges. Still using FlowStats, when the competing WSS is 10MB, its throughput loses $\sim$30for $[1,20K]$ flows, but remains almost unchanged for $[40,60K]$ flows as in Figure 6(a). This is because FlowStats’s hash table grows as a result of growing flow count. Before it fully occupies the last level cache (LLC), increasing the hash table size leads to higher cache miss ratio, which slows down read/write operations and thus throughput. After the LLC is saturated, cache miss ratio stays at a fixed level, causing NFs to exhibit constant throughput. These characteristics also hold in general across NFs as we empirically observe for other NFs in Table 1. This is because generally, (1) NFs typically process either packet headers or payloads, which only depends on several traffic attributes; and (2) traffic attributes usually affect performance by changing the size of key data structures in the NF processing logic, e.g., the mapping table in NAT (slomo20, ; click2000, ; bubbleup11, ), which exhibits the same LLC effect as explained just now.
@@
As shown in Algorithm 1, we first test whether NF performance is sensitive to a traffic attribute or not. Suppose the possible range of an attribute $f$ is $[f_{min},f_{max}]$. We profile the NF’s solo throughput with $f$ set to $f_{min}$ and $f_{max}$ respectively while other attributes remain default (lines 8-9). Then we compare the throughput difference to determine if $f$ should be added in our model. Note that the function profile_one() collects throughput data under a specified configuration (contention level and traffic attribute), and increments the total number of collected samples by one if the configuration has not been profiled. After pruning the attribute list, we carry out a binary search to adaptively collect profiling data. For each call to range_profile(), we consider the difference between solo throughputs on traffic attribute boundaries. If their difference exceeds a certain threshold, we collect $m$ data points at the center of current traffic attribute region (lines 18-22), Then the traffic attribute region is split in half, which will serve as the new region for the recursive call of range_profile().
@@
2:$\mathcal{C}$: The list of performance counters
$\mathcal{F}$: The list of traffic attributes
@@
5:function adaptive_profile($nf$, $\mathcal{C}$, $\mathcal{F}$)
@@
7:    for $f$ in $\mathcal{F}$ do
8:         $T_{min}$leftarrow profile_one($nf$, 00, $f_{min}$, $n$)
9:         $T_{max}$leftarrow profile_one($NF$, 00, $f_{max}$, $n$) $\triangleright$00” represents no contention. $f_{min}$ and $f_{max}$ represent the lowest and highest possible value of $f$ respectively.
@@
11:             $\mathcal{F}\leftarrow\mathcal{F}-\{f\}$triangleright Prune the traffic attribute.
12:    range_profile($nf$, $\mathcal{F}_{min}$, $\mathcal{F}_{max}$, $n$) $\triangleright$ $\mathcal{F}_{min}$ and $\mathcal{F}_{max}$ mean all traffic attributes take the lowest and highest possible value respectively.
13:    return
14:function range_profile($nf$, $\mathcal{F}_{min}$, $\mathcal{F}_{max}$, $n$)
15:    $T_{min}$leftarrow profile_one($nf$, 0, $\mathcal{F}_{min}$, $n$)
16:    $T_{max}$leftarrow profile_one($nf$, 0, $\mathcal{F}_{max}$, $n$)
17:    if $n\geq q$ then $\triangleright$ We reach profiling quota.
18:         return
19:    if $|T_{max}-T_{min}|$geq $\epsilon_{1}$ then
20:         $\mathcal{F}_{mid}$leftarrow $\frac{\mathcal{F}_{max}+\mathcal{F}_{min}}{2}$
21:         for _ in $m$ do
22:             $\mathcal{C}_{r}$leftarrow random() $\triangleright$ Choose a random contention level to apply on target NF.
23:             profile_one($nf$, $\mathcal{C}_{r}$, $\mathcal{F}_{mid}$, $n$)
24:         range_profile($nf$, $\mathcal{F}_{mid}$, $\mathcal{F}_{max}$, $n$)
25:         range_profile($nf$, $\mathcal{F}_{min}$, $\mathcal{F}_{mid}$, $n$)
26:    return
6. Implementation
We implement Yala with $\sim$1600 LoCs in C and Python. Yala employs perf-tools (perf, ) and a working set size estimation tool (wss, ) to collect performance counters. Yala uses sklearn (sklearn, ) to construct machine learning models used in per-resource models. Our code is open source anonymously at (yalagithub, ).
Synthetic benchmarking NFs. We implement three synthetic NFs called mem-bench, regex-bench and compression-bench ($\sim$8300 LoCs in C with DPDK support) based on open-source benchmark tools (stressng, ; mbw, ; RXPbench, ; DPDK, ) to apply configurable levels of contention on memory subsystem, regex, and compression accelerators, respectively.
Network functions. We implement common on-NIC NFs with $\sim$3600 LoCs in C and Click using frameworks including Click 2.1 (click2000, ), DPDK 20.11.6 (dpdk_pl, ), and DOCA 1.5-LTS (doca1, ). Our BF-2 enables hardware flow table offloading for NFs.
7. Evaluation
We present our evaluation of Yala now. The highlights are: